Calculate the change in pH when 7.00 mL of .100 M HCL (aq) is added to 100.0 mL of a buffer solution that is .100 M in NH3 (aq) and .100 M in NH4CL. plse help
Answer #1

pH of a buffer solution is calculated by Henderson equation:

pH = pKa + log[salt]/[acid] ..... for a acid buffer

pOH = pKb + log[salt]/[base] ..... for a basic buffer

This is a basic buffer, so second equation is to be applied here.

When HCl is added to the solution, some NH3 is converted to NH4Cl. As a result concentration of ammonia decreases and concentration of the salt increases.

NH3 + HCl = NH4Cl

Milimoles of HCl = 7 * 0.1 = 0.7

Changed concentration of NH3 = 100*0.1 - 0.7 = 9.3M

Changed concentration of NH4Cl = 100 * 0.1 +0.7 = 10.7

The pKb for ammonia is 4.74

pOH = pKb + log[salt]/[base]

= 4.74 + log10.7/9.3

=4.80

pH = 14 - 4.8

= 9.20